Hello.
School Plays: At a school play, 140 tickets were sold, with total receipts of $290 If adult tickets cost $2.50 each and student tickets cost $1.50 each, how many adult tickets were sold?
Thanks
H.C.

Answer by Zef(4)

Im not entirely sure but the answer should be
1.50 + 2.50(140-x)= 290
1.50 + 350-2.50x=290
351.50-x=290/-2.50
351.50-x= 116
-x=116-351.50
-x= -235.50
x= 235.50
hope i was of help